# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### 26 Jan 2024 v3

@@ 10. Elementary school survey July 1919 2
Next, I overview the survey items and methodology. The investigation was primarily divided into “documentary” and “field” surveys. The former was an analysis of the characteristics of Tsukishima, using published statistics such as the census and the Police Department Statistics Book (Department of Health, Ministry of the Interior 1923a, p. 10). This was a pioneering attempt at the time and had influenced subsequent urban studies (Kawai 1980, pp. 61-62). The latter was a field survey including 14 survey items (Table B.1). It is not possible to describe all the survey items in detail, but following the report format described below, they can be divided into four large categories: 1. survey on workers’ livelihood (chapter 2); 2. survey on sanitary conditions (chapter 3); 3. survey on labor conditions in Tsukishima (chapter 4); and 4. social mapping and photography (appendix). The survey was conducted over two years, generally from November 1918 to November 1920, although the timing varied depending on the survey items (Table B.1).
A unique feature of the survey methodology is using young researchers as commissioned researchers. The household budget survey and leisure life survey were conducted by Gonda Yasunosuke, a faculty member of the Association for German Sciences “Doitsugaku kyōkai gakkō”, who had researched activity photography and entertainment. Tetsuo Hoshino, a medical doctor, was responsible for the sanitary survey. The labor survey was handled by Yoshitsuru Yamana, a graduate of Tokyo Imperial University who contributed to the organization of labor unions. In addition, Toyotaro Miyoshi, a sociology student at the Tokyo Imperial University, and others became temporary employees. In other words, although the Tsukishima Survey was initially planned as an insurance and sanitation survey by the Ministry of Home Affairs, in reality, it was a long-term survey conducted as a joint study by young researchers in different fields, led by Takano (Kawai 1980, p. 63). At the time, the Health Bureau of the Ministry of the Interior was conducting several sanitary surveys of rural areas. Takano might have designed the urban sanitary survey based on this background (Kawai 1981a, pp. 8–9). However, based on the idea that “health and sanitation surveys can at least encompass economic and social surveys,” Takano included a wide range of survey items, not limited to sanitation surveys (Department of Health, Ministry of the Interior 1923a, p. 56). Kawai (1980, p. 55) credits the Tsukishima Survey as a noteworthy pioneering concept: “going beyond a mere insurance and sanitation survey, it attempted to accurately depict the state of urban workers’ lives as they became established in the city during modernization, with skilled workers as its core.”
@@
Takano stated, “To be as complete as possible, it is best to continue the survey over as long a period as possible and to obtain accurate facts by recording the workers’ household income and expenditures over a long period.” For this reason, he believed that “the best method of investigation would be to establish a survey station directly in the survey area and to have a full-time investigator reside there as much as possible, familiarize himself with the area and the people, and conduct the field survey” (Department of Health, Ministry of the Interior 1923a, p. 4). In fact, through the mediation of the Tsukishima police station, a research station was rented at 9-3 Higashinakadori, Tsukishima, where Yoshitsuru Yamana moved in from November 1918 and conducted contact with worker households (Department of Health, Ministry of the Interior 1923a, p. 7; Takeda 2015, p. 77). The survey office was open from November 1918 to December 1920, including the entire survey period (Figure B.1 shows the survey office).
It is also important to note that Takano and his team held weekly meetings at the research center during the survey period. Miyoshi recalls the situation at that time as follows:
After the start of the research station, all the members gathered once a week with Iwasaburo Takano present to discuss the implementation and progress of the survey and administrative matters. Iwasaburo Takano would always attend each meeting, happily listening to the results of the week’s research, inspecting the materials, providing detailed and thoughtful guidance, and usually ending the study session in a friendly atmosphere. Even now, I can recall the warm face of sensei (i.e., Professor Takano) of those days. It was always a friendly, laboratory-like atmosphere in which the C. Booth and S. Rowntree investigations were occasionally discussed.
@@
Kawai (1981b, p. 28) assessed that there were no interconnections among the survey items mentioned above and that the survey “ended up being several fragmentary surveys rather than a comprehensive and systematic survey of workers’ lives.” The Tsukishima Survey is complex for secondary analysis because the sample size is small, and the statistics obtained through the study are not systematically compiled. This difficulty may be why previous studies have not frequently used the Tsukishima Survey. However, as already explained, Takano did not intend to conduct a large-scale social survey in Tsukishima in the first place. Instead, as he noted, the best method is to investigate a representative area “narrowly but deeply,” the goal was to improve the quality of the investigation rather than the scale of it. In this light, the fact that the researcher lived in the survey area and was constantly organizing and revising the survey results should be appreciated. Even though the Tsukishima Survey is the early social survey in Japan, it is of high quality as a prewar household survey.
@@
Finally, I summarize the format of the report of the findings and the unpublished materials this paper will use. The results of the Tsukishima Survey were officially summarized in one report published in December 1921 by the Sanitary Bureau of the Ministry of the Interior and in two appendices (Figure B.2 is an example). The report was published under the name “Report on Field Survey in Tsukishima, Kyobashi, Tokyo,” and consists of four sections. The first section is a comprehensive review by Takano and outlines the survey. The second section summarizes the living conditions of the workers and was written by Gonda (partly by Yamana). Hoshino wrote the third section and summarized the sanitary conditions of workers. In the fourth section, Yamana summarizes the labor situation in Tsukishima. As for the two appendices, one contains statistical tables related to the above report, and the other organizes social maps and photographs. The correspondence between the survey items and the report is summarized in Table B.1. The Ministry of Home Affairs has already published these materials, which are available at the Gonda Yasunosuke Library of the OISR and the University of Tokyo Library.787878 In the OISR, these are included in the Archives of Gonda Yasunosuke (7-2; 7-3; 7-4). In addition, a second manuscript of the Tsukishima Survey report exists. This is the manuscript of the second series, which was preserved in the home of Gonda Hayao, the second son of Gonda Yasunosuke, and was discovered by Terade Koji. The second series includes a summary of the results of a household survey of elementary school teachers’ households in Tokyo, conducted as a supplementary survey to the Tsukishima Survey (see Terade (1982) for details). The manuscript of this second series was later donated to the OISR and is now available as part of the Gonda Yasunosuke Collection at the Institute (The OISR, Archives of Gonda Yasunosuke (8-12; 8-13; 8-14; 8-15)).
@@ B.2 Original Micro-spreadsheets from the Household Budget Survey in Tsukishima Survey
To analyze the consumption behavior of working households in detail, it is necessary to use the original individual survey forms. Unfortunately, the official report does not include individual data on the survey subjects, and many of the original forms used to prepare the report have yet to survive. However, some of the original slips are kept in the archives of the OISR because Takano was the first director of the OISR. The data used in this study to analyze household consumption are the individual data from the household survey (No. 9 in Table B.1), which Gonda used in writing the second section of the official report. While they have yet to be publicly available, I received permission from the OISR to use the original forms stored in the archives: the OISR, Archives of the Tsukishima Survey (THBS, unreleased). [^79] I would like to express my sincere thanks to Professor Kazue Enoki, Deputy Director of the OISR, for allowing me access to the documents. I wish to thank Ms. Mika Nakamura, an archivist at the institute, for her careful assistance during the research.
Figure B.3: An Example Page of the Account of Household Expenses in the Tsukishima Household Budget Survey
@@
Figure B.3: An Example Page of the Account of Household Expenses in the Tsukishima Household Budget Survey
The household budget book used for the household budget survey in the Tsukishima Survey is named the Kinsendeiri hikaechō (money receipts and disbursements account book). This budget book is designed based on the form used in Takano’s first household survey (Online Appendix B.1) (Department of Health, Ministry of the Interior 1923a, p. 152). The budget book is to be filled out daily and includes four sections: income, expenses, receipts from credit purchases, and receipts and disbursements of actual goods. Most of the entries are in the income and expenses sections, indicating that living was basically conducted through wage receipt and consumption in cash. Figure B.3 shows an example of the budget book.
The method used to select households for the survey was not recorded in detail. According to the report, they initially recruited just over 90 households through the factory, police station, and elementary school in Tsukishima. Of these, 50 households completed the book for at least one month, but 10 were deemed incomplete and not included in the report. In addition to the 40 budget books used in preparing the report, therefore, the OISR has a set of other budget books marked as “discarded”. According to my survey, about six incomplete books of the worker households have been determined to be in this “discarded” category. Therefore, based on the remaining conditions of the books, it is plausible to believe that roughly 50 households participated in the survey. Of the 40 households included in the report tally, two households had recorded for more than one year, 11 households had recorded for 6–12 months, and the remaining 27 households had registered for less than six months (Department of Health, Ministry of the Interior 1923a, pp. 152– 153). Initially, the survey period was targeted to be six months, but the average bookkeeping period for the 40 households was four months.
@@ B.3 Trimming
In order to check whether the statistics from 40 THBS budget books contained any missing information, I digitized all 40 household budget books. Throughout this digitization process, I found that there are a few misreporting issues in the budget books.
First, there is a household that does not include information on income (household #37). Although the official report by Gonda reported the average monthly income for this household (Department of Health, Ministry of the Interior 1923a, p. 183), there was no documentation that would allow me to know how this average income was calculated. In order to avoid measurement errors in the income data, I prefer to exclude this household’s budget book.
@@
Thrid, several households did not complete their budget books in the first and/or last month of the survey period. For example, the budget book of household #1 has statistics from 1st January 1919 to 12th January 1920. Similarly, the book of household #29 contains statistics from 26th January to 23rd February 1919. In semi-monthly and monthly datasets, these censored budget books may not be included in the aggregation. In the latter case, for instance, the budget sheets of 26th January and after 11th February 1919 are not included in a semi-month cell because the semi-month cell is defined from the 27th to 11th of the next month in this case. Thus, household #29 only has a cross-sectional observation and can not be included in the panel analysis.
Figure B.4 summarizes the structure of the 40 THBS households. Figure B.4aB.4b, and B.4c shows the observations in the daily, semi-monthly, and adjusted monthly datasets, respectively. As explained, there are five cross-sectional units (#29; #31-33; #35) in the semi-monthly and adjusted monthly datasets. Therefore, there are 35 units with panel structure. After excluding a unit that does not have complete information on income (#37), I get the 34 THBS households as the analytical sample for my panel data analysis.
@@ B.4 Testing the Potential Influence of the Unbalancedness
Size -0.239 0.259 -0.100
[0.251] [0.257] [0.290]
Children aged 6–9 (%) 0.027 -0.003 0.025
[0.027] [0.024] [0.027]
Children aged 10–12 (%) -0.007 -0.029 0.027
[0.029] [0.029] [0.031]
Children aged 13–16 (%) 0.035 -0.045 0.057
[0.039] [0.047] [0.042]
Men aged 17+ (%) -0.002 0.014 -0.008
[0.020] [0.019] [0.026]
Intercept 0.687 -1.142 -0.441
[2.144] [2.042] [2.791]
Wald $\chi^{2}$ statistics for zero slope ($p$-value) 2.03 (0.8444) 2.72 (0.7431) 6.43 (0.2663)
Maximized Log-likelihood -22.20 -19.31 -15.49
Pseudo $R$-squared 0.055 0.063 0.165
Number of households 34 34 34
The results from Probit models are reported. Robust standard errors are in brackets. Notes: The dependent variable is an indicator variable that takes one if the household has shorter time-series observations. Column 1 uses the median of the number of days per unit as the threshold. The result from the regression using the average number of days per unit ($146$ days) as the threshold is identical to that listed in column 1. Column 2 uses the 75 percentile of the number of days per unit as the threshold. Column 3 uses the 25 percentile of the number of days per unit as the threshold. All family size variables are the time-series average in each unit over the sample period. The proportion of children aged 0–5 years (%) is used as the reference group.
The unbalancedness in the THBS sample suggests that the attrition in several units may lead to selection bias in the sample. If the difference in the preference for consumptions predicted the large attritions, there must be statistically significant correlations between the family size variables and attritions. In other words, the units with shorter (longer) time-series observations have different preferences from those with longer (shorter). Table B.2 summarizes the results of the balancing tests for the 34 THBS households used in the analyses (Section 3.2). The dependent variable is an indicator variable that takes one if the household has time-series observations less than the median (column 1), 75 percentile (column 2), and 25 percentile (column 3). In Column 1, all the estimated coefficients on the family size variables are very close to zero and statistically insignificant, and the Wald statistics suggest the null results. This result is unchanged if I use the third and first quantiles in columns (2) and (3), respectively. Therefore, the family characteristics are similar in a statistical sense between the households with short-term observations and those with long-term observations. This supports the evidence that the unbalancedness in my THBS household panel datasets is not problematic in estimating their consumption behaviors.

### 25 Feb 2024 v6

@@ 10. Elementary school survey July 1919 2
Next, I overview the survey items and methodology. The investigation was primarily divided into “documentary” and “field” surveys. The former was an analysis of the characteristics of Tsukishima, using published statistics such as the Police Department Statistics Book (Department of Health, Ministry of the Interior 1923a, p. 10). This was a pioneering attempt at the time and had influenced subsequent urban studies (Kawai 1980, pp. 61-62). The latter was a field survey including 14 survey items (Table B.1). It is not possible to describe all the survey items in detail, but following the report, they can be divided into four large categories: 1. survey on workers’ livelihood (chapter 2); 2. survey on sanitary conditions (chapter 3); 3. survey on labor conditions (chapter 4); and 4. social mapping and photography (appendix). The survey was conducted over two years, generally from November 1918 to November 1920, although the timing varied depending on the survey items (Table B.1).
A unique feature of the survey methodology is using young researchers as commissioned researchers. The household budget survey and leisure life survey were conducted by Yasunosuke Gonda, a faculty member of the Doitsugaku kyōkai gakkō (Association for German Sciences), who had researched activity photography and entertainment. Tetsuo Hoshino, a medical doctor, was responsible for the sanitary survey. The labor survey was handled by Yoshitsuru Yamana, a graduate of Tokyo Imperial University who contributed to the organization of labor unions. In addition, Toyotaro Miyoshi, a sociology student at the Tokyo Imperial University, and others became temporary employees. In other words, although the Tsukishima Survey was initially planned as an insurance and sanitation survey by the Ministry of Home Affairs, in reality, it was a long-term survey conducted as a joint study by young researchers in different fields, led by Takano (Kawai 1980, p. 63). At the time, the Health Bureau of the Ministry of the Interior was conducting several sanitary surveys of rural areas. Takano may have designed the urban sanitary survey from this background (Kawai 1981a, pp. 8–9). However, based on the idea that “health and sanitation surveys can at least encompass economic and social surveys,” Takano included a wide range of survey items, not limited to sanitation surveys (Department of Health, Ministry of the Interior 1923a, p. 56). Kawai (1980, p. 55) credits the Tsukishima Survey as a noteworthy pioneering concept: “going beyond a mere insurance and sanitation survey, it attempted to accurately depict the state of urban workers’ lives as they became established in the city during modernization, with skilled workers as its core.”
@@
Takano stated, “To be as complete as possible, it is best to continue the survey over as long a period as possible and to obtain accurate facts by recording the workers’ household income and expenditures over a long period.” For this reason, he believed that “the best method of investigation would be to establish a survey station directly in the survey area and to have a full-time investigator reside there as much as possible, familiarize himself with the area and the people, and conduct the field survey” (Department of Health, Ministry of the Interior 1923a, p. 4). In fact, through the mediation of the Tsukishima police station, a survey office was rented at 9-3 Higashinakadori, Tsukishima, where Yoshitsuru Yamana moved in from November 1918 and contacted worker households (Department of Health, Ministry of the Interior 1923a, p. 7; Takeda 2015, p. 77). The survey office was open from November 1918 to December 1920, including the entire survey period. Figure B.1 shows the photographs of the survey office.
It is also important to note that Takano and his team held weekly meetings at the research center during the survey period. Toyotarō Miyoshi, who was an assistant in the survey, recalls the situation at that time as follows:
After the start of the survey office, all the members gathered once a week with Iwasaburo Takano present to discuss the implementation and progress of the survey and administrative matters. Iwasaburo Takano would always attend each meeting, happily listening to the results of the week’s research, inspecting the materials, providing detailed and thoughtful guidance, and usually ending the study session in a friendly atmosphere. Even now, I can recall the warm face of sensei (i.e., Professor Takano) of those days. It was always a friendly, laboratory-like atmosphere in which the C. Booth and S. Rowntree investigations were occasionally discussed.
@@
Kawai (1981b, p. 28) assessed that there were no interconnections among the survey items mentioned above and that the survey “ended up being several fragmentary surveys rather than a comprehensive and systematic survey of workers’ lives.” This difficulty may be why previous studies have not frequently used the Tsukishima Survey. However, as already explained, Takano did not intend to conduct a large-scale social survey in Tsukishima in the first place. Instead, as he noted, the best method is to investigate a representative area “narrowly but deeply,” the goal was to improve the quality of the investigation rather than the scale of it. In this light, the fact that the researcher lived in the survey area and was constantly organizing and revising the survey results should be appreciated. Even though the Tsukishima Survey is the early social survey in Japan, it is of high quality as a prewar household survey.
@@
Finally, I summarize the format of the published report and the unpublished materials. The results of the Tsukishima Survey were officially summarized in one report published in December 1921 by the Sanitary Bureau of the Ministry of the Interior and in two appendices (Figure B.2 is an example of the front cover). The report was published under the name “Report on Field Survey in Tsukishima, Kyobashi, Tokyo,” and consists of four sections. The first section is a comprehensive review by Takano and outlines the survey. The second section summarizes the living conditions of the workers and was written by Gonda (partly by Yamana). Hoshino wrote the third section and summarized the sanitary conditions of workers. In the fourth section, Yamana summarizes the labor situation in Tsukishima. As for the two appendices, one contains statistical tables related to the above report, and the other organizes social maps and photographs. The correspondence between the survey items and the report is summarized in Table B.1. The Ministry of Home Affairs has already published these materials, which are available at the Gonda Yasunosuke Library of the OISR and the University of Tokyo Library.848484 In the OISR, these are included in the Archives of Gonda Yasunosuke (7-2; 7-3; 7-4). In addition, a second manuscript of the Tsukishima Survey report exists. This is the manuscript of the second series, which was preserved in the home of Gonda Hayao, the second son of Gonda Yasunosuke, and was discovered by Terade Koji. The second series includes a summary of the results of a household survey of elementary school teachers’ households in Tokyo, conducted as a supplementary survey to the Tsukishima Survey (see Terade (1982) for details). The manuscript of this second series was later donated to the OISR and is now available as part of the Gonda Yasunosuke Collection at the Institute (The OISR, Archives of Gonda Yasunosuke (8-12; 8-13; 8-14; 8-15)).
@@ B.2 Original Micro-spreadsheets from the Household Budget Survey in Tsukishima Survey
To analyze the consumption smoothing behavior of THBS households, it is necessary to use the original micro-spreadsheets. The official report does not include household-level information, and a large part of the original forms used to prepare the report have little survived. Fortunately, however, some of the original micro-spreadsheets are kept in the archives of the OISR because Takano was its first director. This study uses the original micro-spreadsheets from the household budget survey (No. 9 in Table B.1), which was used in writing the second section of the official report by Gonda. While they have yet to be publicly available, I received permission from the OISR to use the original forms stored in the archives: the OISR, Archives of the Tsukishima Survey (THBS, unreleased). [^85] I would like to express my sincere thanks to Professor Kazue Enoki, Deputy Director of the OISR, for allowing me access to the documents. I wish to thank Ms. Mika Nakamura, an archivist at the institute, for her careful assistance during the research.
Figure B.3: An Example Page of the Budget Book in the Tsukishima Household Budget Survey
@@
Figure B.3: An Example Page of the Budget Book in the Tsukishima Household Budget Survey
The budget book used for the household budget survey is named the Kinsendeiri hikaechō (money receipts and disbursements account book). This budget book is designed based on the form used in Takano’s first household survey (Online Appendix B.1). [^86] Department of Health, Ministry of the Interior (1923a, p. 152). The budget book is to be filled out daily and includes four sections: income, expenses, receipts from credit purchases, and receipts and disbursements of actual goods. Most of the entries are in the income and expenses sections, indicating that living was basically conducted through wage receipt and consumption in cash. Figure B.3 shows an example of the budget book.
As described, the method used to select households for the survey was not recorded in detail (Section B.1). According to the report, they initially recruited just over 90 households through the factory, police station, and elementary school in Tsukishima. Of these, 50 households completed the book for at least one month, but 10 were deemed incomplete and not included in the report. In addition to the 40 budget books used in preparing the report, therefore, the OISR has a set of other budget books marked as “discarded”. According to my survey, about six incomplete books of the worker households have been determined to be in this “discarded” category. Therefore, based on the remaining conditions of the books, it is plausible to believe that roughly 50 households participated in the survey. Of the 40 households included in the report tally, two households had recorded for more than one year, 11 households had recorded for 6–12 months, and the remaining 27 households had registered for less than six months (Department of Health, Ministry of the Interior 1923a, pp. 152– 153). Initially, the survey period was targeted to be six months, but the average bookkeeping period for the 40 households was four months.
@@ B.3 Trimming
To check whether the statistics from THBS budget books contained any missing information, I digitized all 40 household budget books. Throughout this digitization process, I found that there are a few misreporting issues in the budget books.
First, there is a household that does not include income information (household #37). Although the official report by Gonda reported the average monthly income for this household (Department of Health, Ministry of the Interior 1923a, p. 183), no documentation would allow me to know how this average income was calculated. Hence, I excluded this household’s budget book. Similarly, a household does not include information on the payments to the credit purchases (household #39). Although this household’s budget book documented the amount purchased in credit purchases, I excluded this household’s budget book because the expenditure series becomes systematically smaller due to the omission of the payment information.
@@
Third, several households did not complete their budget books in the first and/or last month of the survey period. For example, the budget book of household #1 has statistics from 1st January 1919 to 12th January 1920. Similarly, the book of household #29 contains statistics from 26th January to 23rd February 1919. In semi-monthly and monthly datasets, these censored budget books may not be included in the aggregation. In the latter case, for instance, the budget sheets of 26th January and after 11th February 1919 are not included in a semi-month cell because the semi-month cell is defined from the 27th to 11th of the next month in this case. Thus, household #29 only has a cross-sectional observation and can not be included in my analytical sample.
Figure B.4 summarizes the structure of the 40 THBS households. Figure B.4aB.4b, and B.4c shows the observations in the daily, semi-monthly, and adjusted monthly datasets, respectively. As explained, there are five cross-sectional units (#29; #31-33; #35) in the semi-monthly and adjusted monthly datasets. Therefore, there are 35 units with a panel structure. After excluding a unit without complete income information (#37), I get the 34 THBS households as the analytical sample for my panel data analysis.
@@ B.4 Testing the Potential Influence of the Unbalancedness
Size -0.243 0.238 -0.135
[0.244] [0.254] [0.294]
Children aged 6–12 (%) 0.017 -0.011 0.026
[0.020] [0.019] [0.022]
Children aged 13–16 (%) 0.029 -0.050 0.057
[0.038] [0.045] [0.041]
Men aged 17+ (%) -0.008 0.008 -0.010
[0.019] [0.018] [0.026]
Intercept 1.149 -0.576 -0.164
[2.107] [2.048] [2.805]
Zero slope ($p$-value) 0.710 0.753 0.160
Maximized Log-likelihood -21.8 -18.5 -15.2
Pseudo $R$-squared 0.046 0.046 0.171
Number of households 33 33 33
The results from Probit models are reported. Robust standard errors are in brackets. Notes: The dependent variable is an indicator variable that takes one if the household has shorter time-series observations. Column 1 uses the median of the number of days per unit as the threshold. The result from the regression using the average number of days per unit ($144$ days) as the threshold is identical to that listed in column 1. Column 2 uses the 75 percentile of the number of days per unit as the threshold. Column 3 uses the 25 percentile of the number of days per unit as the threshold. All family size variables are the time-series average in each unit over the sample period. The proportion of children aged 0–5 years (%) is used as the reference group. Wald $\chi^{2}$ statistics $p$-values for the null of the zero-slope hypothesis are reported in the sixth row. The results from the expanded regressions including the head’s age and earnings are materially similar. The null of the zero-slope hypothesis is not rejected in all the specifications at the conventional level (not reported).
In this subsection, I test whether attrition leads to selection bias in the THBS sample. If the difference in the preference for consumptions predicted the large attritions, there must be statistically significant correlations between the family size variables and attritions. In other words, the units with shorter (longer) time-series observations have different preferences from those with longer (shorter) time-series observations. Table B.2 summarizes the results of the balancing tests (Section 3.2). The dependent variable is an indicator variable that takes one if the household has time-series observations less than the median (column 1), 75 percentile (column 2), and 25 percentile (column 3). In Column 1, all the estimated coefficients on the family size variables are close to zero and statistically insignificant, and the Wald statistics $p$-value suggests the null results. This is unchanged if I use the third and first quantiles in columns (2) and (3), respectively. Therefore, the family characteristics are similar between the households with short-term observations and those with long-term observations. This result supports the evidence that the unbalancedness in the THBS dataset is unlikely to lead to selection bias.